Solve for x: 2log(3x)-log(8x+1)=log x I think the answer is x=0,1/2 ????
Answer by edjones(8007)

2log(3x)-log(8x+1)=log x
log((3x)^2/(8x+1))=log x
10^log((3x)^2/(8x+1))=10^log x
(3x)^2/(8x+1)=x
(3x)^2=x(8x+1)
9x^2=8x^2+x
x^2-x=0
x(x-1)=0
x=0, x=1
x=0 is not an answer because 2log(3*0) is undefined.
x=1 is the only answer.
.
Ed